OpenStudy (maeganr):

Standard Form? 10 Problems..

OpenStudy (maeganr):

"Find the x- and y- intercepts of the graph of each equation." #1. x+y=7 #3. 2x+3y=-6 #5. 5x-4y=-12 "Draw a line with the given intercepts." #7. x-intercept: 4 y-intercept: 5 #9. x-intercept: -6 y-intercept: -8 "Graph each equation using x- and y- intercepts." #11. -3x-6y=12 "For each equation, tell whether its graph is a horizontal or vertical line." #13. y=-2 #15. y=-0.25 "Graph each equation." #17. y=6 #19. y=-7 EVEN IF YOU HELP ME WITH ONE OF EACH FROM EACH SET OF INSTRUCTIONS THAT WOULD BE GREAT. THANKS FOR ALL OF YOU WHO HELP ME! (:

OpenStudy (anonymous):

for finding the x and y intercepts just cover the other variable .. for example x+y= 7 cover the x and you get y=7 cover the y and you get x= 7

OpenStudy (maeganr):

so its 7?

OpenStudy (anonymous):

yup, I believe so

OpenStudy (anonymous):

try the same thing with the next two, when you think you have the answer, tell me it and I will tell u if its right

OpenStudy (anonymous):

ok?

OpenStudy (maeganr):

crap. i just did #2 not #3 i only have to do odd. srry brb

OpenStudy (anonymous):

kk- r u using a textbook?

OpenStudy (maeganr):

y=-2 x=-3

OpenStudy (anonymous):

yup ☺

OpenStudy (maeganr):

we do have textbooks but my teacher doesn't let us use it. we only have what frayed diagrams she gives us and our workbook with just problems in it. ): the textbook would be reeeeally useful.

OpenStudy (anonymous):

oh..

OpenStudy (maeganr):

yeah it sucks

OpenStudy (anonymous):

yup

OpenStudy (maeganr):

#5. y=-3 x= -2.4 ?

OpenStudy (anonymous):

y= 3 but yah

OpenStudy (maeganr):

oh yeah i see now #7 how do i do it. and btw i was out sick the whole week she taught all this so i never learned it so you're teaching me right now. (:

OpenStudy (anonymous):

Oh.. ok ☺ I feel your pain, I was sick and I had to teach myself

OpenStudy (maeganr):



OpenStudy (maeganr):

#7. ?